Construct a​ 99% confidence interval to estimate the population mean using the data below.    

x overbar=21 s=5.5 n=25 N=180
Answer #1

Date: Solutions - Given that x=21 s=5.5 n=25 dFanal = 25-1 = 24 At 99% confidence level the tis L=1-99 % do 1-0.99 200 a oot